# Simulating discrete-time quantum walk with urn model

### v1

@@ Abstract
Urn models have long been used to study stochastic processes, probability distributions, and reinforcement dynamics. Meanwhile, discrete time quantum walks (DTQW) serve as fundamental components in quantum computation and quantum information theory. This work explores a novel connection between an urn model and discrete-time quantum walks, focusing on how urn-based stochastic processes can provide insights into quantum state evolution and algorithmic behavior. Importantly, our model operates entirely with real-valued quantities, without relying on complex amplitudes of the quantum walk. Every stochastic rule is derived from urn-based processes, whose parameters are governed by the coin and shift mechanisms of the quantum walk, maintaining the structural dynamics of the QW. Through this interdisciplinary approach, we aim to bridge classical and quantum frameworks, offering new perspectives on both quantum algorithms and stochastic modeling. This work connects the urn model to the broad range of applications where DTQWs are successfully employed.
Exploring the dynamics of a physical system often involves identifying suitably engineered dynamics in a potentially very different realm that exhibits analogous behavior. This approach not only enhances our understanding of the physical system but also enriches the dynamics of that different realm, making both perspectives more insightful and beneficial. A classic example emerged in 2002, when Grover’s quantum search algorithm was shown to correspond to the behavior of coupled pendulums [1]. Recent work [2] clarified this connection further. It showed that the dynamics of any classical system of harmonic oscillators can be equivalently understood as the dynamics of a corresponding quantum system of exponentially smaller size. Conversely, every quantum algorithm can be equivalently understood as the propagation of a classical wave in a system of coupled oscillators, albeit one that may involve action at a distance and significant classical resources, which make the latter process impractical. But the important point is that such an analogy led to the discovery of an efficient quantum algorithm to simulate the classical dynamics of coupled oscillators.
This duality underscores the importance of systematically studying how useful classical systems—such as coupled oscillators, which can themselves simulate a wide variety of classical dynamics—can reflect the behavior of quantum algorithms. While such classical simulations might incur substantial computational costs or non-local characteristics, they can lead to the development of novel quantum algorithms for simulating intricate classical systems, as well as quantum-inspired classical strategies. In this spirit, we employ discrete-time quantum walks (DTQW) and construct an urn model as the corresponding useful classical system, offering a fresh route to replicate quantum dynamics.
@@
Quantum walks on graphs are one of the few established design strategies for quantum algorithms. Despite their simple description, quantum walks are extremely powerful, as they are universal for quantum computation [3, 4]. Owing to their non-classical spreading dynamics, localization property, and potential algorithmic advantages, quantum walks have been extensively explored both theoretically and experimentally [5, 6, 7].
On the other hand, Urn models—built on the simple act of drawing and replacing colored balls—serve as a versatile framework for modeling both random and self‑reinforcing dynamics across domains as varied as population genetics, network growth, statistical physics, and machine learning [8, 9].
@@
Replicating quantum phenomena like DTQWs on useful classical platforms is inherently challenging. This difficulty stems from the fact that DTQWs rely on quantum superposition, where evolving states interfere—constructively or destructively—based on their relative phases. As a result, they produce probability distributions that differ from those of classical random walks.
@@
A line of research on this topic has approached this challenge by constructing classical random walks (CRWs) [10, 11] whose space and time-dependent transition probabilities in each step are extracted from the complex amplitudes of the DTQW. This approach yields a non-homogeneous CRW whose probability distribution replicates that of the corresponding DTQW. Although effective in reproducing certain statistics, these methods necessarily break the spatial and temporal structure of the DTQW and demand access to quantum amplitudes at each step, an approach analogous to Bohmian mechanics[12], where the quantum wave function guides particle trajectories.
@@
In contrast, alternative formulations of quantum mechanics—akin to the hydrodynamic approaches of Holland and Poirier [13, 14, 15, 16, 17, 18] —eschew direct reliance on the wave function, instead constructing real-valued fields or potentials that reproduce quantum evolution without amplitude-based guidance. Inspired by this distinction, we propose an urn‑model framework for simulating DTQWs that never invokes complex amplitudes. All transition rules arise from real‑valued urn processes whose parameters depend only on the QW’s coin and shift structure. Crucially, our model mirrors the structural properties of the original quantum walk system. When the QW has uniform coin and shift operations, our urn model remains homogeneous in space and time, employing identical urn parameters at each node and time step. If QW coin or shift operations vary across space or time, the corresponding urn model adapts its parameters in the same pattern, preserving the structural variations of QW. This yields two main advantages. First, preserving the structural variations of the QW simplifies both the analysis and the simulation, while allowing our method to be naturally extended to higher-dimensional graphs. Second, by avoiding amplitude-based guidance, the urn model brings into focus a fundamental probabilistic mechanism underlying quantum walk spreading and interference, insights that may carry foundational significance.
@@
$S=\bigl{(}\lvert 0\rangle\langle 0\rvert\otimes\sum_{n=-\infty}^{\infty}\lvert n+1\rangle\langle n\rvert\bigr{)}+\bigl{(}\lvert 1\rangle\langle 1\rvert\otimes\sum_{n=-\infty}^{\infty}\lvert n-1\rangle\langle n\rvert\bigr{)}.$ (2)
@@
The process of $W$ is iterated without resorting to the intermediate measurements to realize a large number of steps of the quantum walk.
@@
is sufficient to describe the most general form of the discrete-time quantum walk, where $\xi,\theta,\zeta\in\mathbb{R}$ and $\theta\in[0,\pi/2]$. By choosing different $(\xi,\theta,\zeta)$, one obtains a variety of discrete-time walks with distinct dynamical features.
@@
Suppose $\psi_{0}^{\prime}=r_{0}^{\prime}e^{i\phi_{0}^{\prime}}$ and $\psi_{1}^{\prime}=r_{1}^{\prime}e^{i\phi_{1}^{\prime}}$ then one can write
$\begin{array}[]{rc}P_{0}^{\prime}={r_{0}^{\prime}}^{2}=\frac{1}{2}(r_{0}^{2}+r_{1}^{2}+2r_{0}r_{1}\cos(\phi_{0}-\phi_{1})),\\ P_{1}^{\prime}={r_{1}^{\prime}}^{2}=\frac{1}{2}(r_{0}^{2}+r_{1}^{2}-2r_{0}r_{1}\cos(\phi_{0}-\phi_{1})),\end{array}$ (12)
and
$\begin{array}[]{rc}\phi_{0}^{\prime}=\tan^{-1}\left(\frac{r_{0}\sin\phi_{0}+r_{1}\sin\phi_{1}}{r_{0}\cos\phi_{0}+r_{1}\cos\phi_{1}}\right),\\ \phi_{1}^{\prime}=\tan^{-1}\left(\frac{r_{0}\sin\phi_{0}-r_{1}\sin\phi_{1}}{r_{0}\cos\phi_{0}-r_{1}\cos\phi_{1}}\right).\end{array}$ (15)
Where $P_{0}^{\prime}$ and $P_{1}^{\prime}$ are resulting probabilities of measuring the transformed state in the computational basis. We now construct an urn model specifically designed to replicate the measurement statistics of a single-qubit state (Eq. 8), as well as the statistics following the application of a Hadamard operation on that state (Eq. 12 and 15).
An urn is a container that holds a collection of objects, usually called balls. The contents of the urn can be altered by randomly removing or adding balls according to the rules given. The behavior of such a system is probabilistic in nature. In the literature, it has been noted earlier that urn models has the capability to describe various quantum logic [20, 21].
Imagine a setting with two urns, labeled 00 and $1$, and a total collection of $N$ balls. The various operations on this system are categorized into three classes: preparation ($\mathbb{P}$), transformation ($\mathbb{T}$), and measurement ($\mathbb{M}$).
@@
In a typical $\mathbb{P}$ procedure, we begin by selecting one ball uniformly at random from the collection of $N$ balls and marking it for future use. Following this, the $N$ balls (including the marked ball) are distributed between the two urns. The allocation can proceed in one of two ways, each chosen with equal probability:
@@
  • Either $N_{0}=\lfloor\rho_{0}N\rfloor$ balls are randomly selected from the collection and placed into urn 00, with the remaining $N_{1}=N-N_{0}$ balls allocated to urn 1;
@@
  • Or $N_{0}=\lceil\rho_{0}N\rceil$ balls are randomly selected from the collection and placed into urn 0, with the remaining $N_{1}=N-N_{0}$ balls placed into urn 1.
Here, the parameter $\rho_{0}\in[0,1]$ governs the expected proportion of balls assigned to urn 00. It is important to note that, during the selection procedure, all balls—including the marked one—are treated as identical. Additionally, each urn $i$ (for $i=0,1$) is assigned a real-valued phase tag $\eta_{i}$. The prepared state is denoted by $\textit{U}^{N,N_{0}}_{\eta_{0},\eta_{1}}$.
Now, let us define the transformation process $\mathbb{T}$ on the state $\textit{U}^{N,N_{0}}_{\eta_{0},\eta_{1}}$. In the $\mathbb{T}$ process, some randomly selected balls from one of the urns are transferred to the other urn according to a certain stochastic rule. As a result, the number of balls in each urn is updated and becomes either $N_{0}^{\prime}=\lfloor\tilde{N}_{0}\rfloor$ and $N_{1}^{\prime}=N-\lfloor\tilde{N}_{0}\rfloor$, or $N_{0}^{\prime}=\lceil\tilde{N}_{0}\rceil$ and $N_{1}^{\prime}=N-\lceil\tilde{N}_{0}\rceil$, each with equal probability. Here, $\tilde{N}_{0}$ is a function of $N_{0},N_{1},\eta_{0}$ and $\eta_{1}$.
The phase tags $\eta_{0}$ and $\eta_{1}$ are also updated to new values $\eta_{0}^{\prime}$ and $\eta_{1}^{\prime}$, where $\eta_{i}^{\prime}=\eta_{i}^{\prime}(N_{0},N_{1},\eta_{0},\eta_{1})$.
Putting it all together, the transformation is described as:
$\mathbb{T}:\textit{U}^{N,N_{0}}_{\eta_{0},\eta_{1}}\longrightarrow\textit{U}^{N,N_{0}^{\prime}}_{\eta_{0}^{\prime},\eta_{1}^{\prime}}$
@@
Now, let us define a process referred to as the measurement process, denoted by $\mathbb{M}$. In the $\mathbb{M}$ process, all balls are placed into the urn that contains the marked ball, and the measurement outcome is the number associated with that urn. Both phase tags are then reset to zero.
Thus, the post-measurement state becomes:
  • $U^{N,N}_{0,0}$ if the marked ball is in urn 0;
  • $U^{N,0}_{0,0}$ if the marked ball is in urn 1.
For example, suppose that after a $\mathbb{T}$ process the marked ball belongs to urn 1. Then, the measurement outcome will be 1, and all $N$ balls will be placed into urn 1. If the $\mathbb{M}$ process is applied again to the urn system, the outcome will remain the same, because the marked ball will always be found in urn 1.
Now, suppose that $\mathbb{P}$ and $\mathbb{M}$ are applied to the system without any $\mathbb{T}$ process in between. In the $\mathbb{P}$ process, we set $\eta_{i}=\phi_{i}$ and $\rho_{0}=r_{0}^{2}$. The probability of randomly selecting a particular ball to be marked from $N$ identical balls is $\frac{1}{N}$. As a result, the probability of obtaining outcomes 0 and 1 after the $\mathbb{M}$ process, for a particular configuration, will be $\frac{N_{0}}{N}$ and $\frac{N_{1}}{N}$, respectively.
Thus, the ensemble average of the probabilities corresponding to outcomes 0 and 1 can be written as:
@@
$P_{0}=\left\langle\frac{N_{0}}{N}\right\rangle=\frac{\lfloor r_{0}^{2}N\rfloor+\lceil r_{0}^{2}N\rceil}{2N}$
and
$P_{1}=1-\left\langle\frac{N_{0}}{N}\right\rangle$
respectively.
Now, one can always write:
$r_{0}^{2}N=\lfloor r_{0}^{2}N\rfloor+\epsilon\quad\text{and}\quad r_{0}^{2}N=\lceil r_{0}^{2}N\rceil-\epsilon^{\prime}$
where $10\leq\epsilon,\epsilon^{\prime}<$ 1. Moreover, whenever $r_{0}^{2}N\notin\mathbb{Z}$, it follows that $\epsilon+\epsilon^{\prime}=1$.
Hence,
$\lfloor r_{0}^{2}N\rfloor+\lceil r_{0}^{2}N\rceil=2r_{0}^{2}N+(\epsilon^{\prime}-\epsilon)$
and thus,
@@
$\lim_{N\to\infty}\frac{\lfloor r_{0}^{2}N\rfloor+\lceil r_{0}^{2}N\rceil}{2N}=\lim_{N\to\infty}\left(r_{0}^{2}+\frac{\epsilon^{\prime}-\epsilon}{2N}\right)=r_{0}^{2}$
Therefore, in the large $N$ limit, we obtain $P_{0}=r_{0}^{2}$ and $P_{1}=1-r_{0}^{2}=r_{1}^{2}$.
Now, let us define a specific transformation denoted by $\mathbb{T}_{H}$ acting on $\textit{U}^{N,N_{0}}_{\phi_{0},\phi_{1}}$. This transformation is constructed such that when followed by a $\mathbb{M}$ operation, it produces measurement statistics that converge to the Hadamard operation (Eq. (12)) in the limit $N\to\infty$. Furthermore, the updated phase tags also converge to $\phi_{0}^{\prime}$ and $\phi_{1}^{\prime}$ (as given by Eq. (15)) in that limit.
@@
So, the new quantity $\tilde{N}_{0}$ in the transformation $\mathbb{T}_{H}$ must be chosen such that, as $N\to\infty$, the probability of obtaining the measurement outcome 00 becomes
@@
and the probability of outcome $1$ becomes
@@
Furthermore, we require
$\lim_{N\to\infty}\langle\eta_{i}^{\prime}(N_{0},N_{1},\phi_{0},\phi_{1})\rangle={\phi_{i}}^{\prime}$ (18)
where $\eta_{i}^{\prime}$ represents the updated phase tags.
To implement this, we adopt the following transformation scheme [see appendix]:
$\tilde{N}_{0}=\frac{N_{0}+N_{1}}{2}+\sqrt{N_{0}N_{1}}\cos(\phi_{1}-\phi_{0}).$ (19)
The updated phase tags are given by
$\begin{array}[]{rc}\eta_{i}^{\prime}(N_{0},N_{1},\phi_{0},\phi_{1})=\tan^{-1}\left(\frac{\sqrt{\frac{N_{0}}{N}}\sin\phi_{0}+(-1)^{i}\sqrt{\frac{N_{1}}{N}}\sin\phi_{1}}{\sqrt{\frac{N_{0}}{N}}\cos\phi_{0}+(-1)^{i}\sqrt{\frac{N_{1}}{N}}\cos\phi_{1}}\right)\end{array}$ (21)
According to this scheme, the updated number of balls in urn 0, denoted by $N_{0}^{\prime}$, will either be $\lfloor\frac{N}{2}+\sqrt{N_{0}N_{1}}\cos(\phi_{1}-\phi_{0})\rfloor$ or $\lceil\frac{N}{2}+\sqrt{N_{0}N_{1}}\cos(\phi_{1}-\phi_{0})\rceil$, each chosen with equal probability. If $N_{0}^{\prime}>N_{0}$, the $\mathbb{T}_{H}$ adds $N_{0}^{\prime}-N_{0}$ balls to urn 0 from urn 1. Conversely, if $N_{0}^{\prime}<N_{0}$, it moves $N_{0}-N_{0}^{\prime}$ balls from urn 0 to urn 1.
One can show that [see appendix] in the limit $N\to\infty$, the measurement outcome probabilities converge to those of the Hadamard transformation.
For finite $N$, there will always be some error, but by taking $N$ large enough, one can approximate such quantum operations with arbitrary precision.
Moreover, this method can be generalized to simulate other single-qubit quantum coin operations, as expressed in Eq. (4). One can obtain the generalized forms of $\tilde{N}_{0}$ and $\eta_{i}^{\prime}$ in the corresponding transformation process denoted by $\mathbb{T}_{c}$ as follows (see appendix):
$\begin{array}[]{rc}\tilde{N}_{0}=N_{0}\cos^{2}\theta+N_{1}\sin^{2}\theta&\\ &+2\sqrt{N_{0}N_{1}}\sin\theta\cos\theta\cos(\phi_{0}-\phi_{1}+\xi-\zeta)\\ \end{array}$ (24)
$\begin{array}[]{rc}\eta_{i}^{\prime}=\tan^{-1}\left(\frac{\sqrt{\frac{N_{0}}{N}}\cos\theta\sin(\xi+\phi_{0})+(-1)^{i}\sqrt{\frac{N_{1}}{N}}\sin\theta\sin(\zeta+\phi_{1})}{\sqrt{\frac{N_{0}}{N}}\cos\theta\cos(\xi+\phi_{0})+(-1)^{i}\sqrt{\frac{N_{1}}{N}}\sin\theta\cos(\zeta+\phi_{1})}\right)\end{array}$ (26)
Finally, we define the urn model corresponding to a DTQW using the generalized $SU(2)$ coin (Eq. (4)), with the initial state given by:
Figure 1: Probability distributions arising from the $\mathbb{M}$ process in the urn model of the Hadamard coined quantum walk, with initial state $\frac{1}{\sqrt{2}}\lvert 0\rangle-\frac{i}{\sqrt{2}}\lvert 1\rangle$ and at time $t=100$, are shown for different total numbers of balls $N$. As $N$ increases, the distribution converges to that of the ideal Hadamard quantum walk.
@@
We perform the $\mathbb{P}$, $\mathbb{T}_{c}$, and $\mathbb{M}$ processes within a generalized framework. We begin with a collection of $N$ balls. To each lattice site $x\in\mathbb{Z}$, we assign two urns, labeled $(x,0)$ and $(x,1)$. One of the $N$ balls is randomly selected and marked, after which all $N$ balls are distributed among the urns.
The state of the system at time $t$ is fully specified by the nonnegative integers
$N_{x,c}(t)\;=\;\text{number of balls in urn }(x,c),\quad c\in\{0,1\},$
together with associated “phase” tags $\eta_{x,c}(t)\in[0,2\pi)$ for each occupied urn. These obey the global constraint
@@
At time $t=0$, only the two urns at the origin carry balls; all others are empty:
@@
Within the origin urns, we set

### v3

# Emergent Quantum Walk Dynamics from Classical Interacting Particles

@@ Abstract
The dynamics of a discrete-time quantum walk (DTQW) can be realized within a purely classical interacting particle system composed of some boxes and a large but finite number of balls, and can, in principle, be implemented in a tabletop experimental setting. The distribution of the balls evolves under stochastic, occupation-dependent update rules at each lattice site, producing quantum-walk dynamics without invoking a wavefunction. The update parameters are fixed by the parameters of coin and shift operations of the DTQW. This framework naturally yields a generalized active spin model and provides a minimal lattice-based microscopic understanding of the emergence of quantum-like dynamics in active matter systems. This interdisciplinary approach connects the classical models to the broad range of applications where DTQWs are successfully employed.
@@
Discrete-time quantum walks (DTQW) on graphs serve as a well-established framework for developing quantum algorithms. Though conceptually straightforward, they possess remarkable computational power and have been shown to be universal for quantum computation [1, 2]. Owing to their non-classical spreading dynamics, localization property, and potential algorithmic advantages, DTQW have been extensively explored both theoretically and experimentally [3, 4, 5].
@@
Modeling quantum phenomena like DTQWs on classical platforms is inherently challenging. This difficulty stems from the fact that DTQWs rely on quantum superposition, where evolving states interfere—constructively or destructively—based on their relative phases. As a result, they produce probability distributions that differ from those of classical random walks.
@@
A line of research on this topic has approached this challenge by constructing classical random walks (CRWs) [6, 7, 8] whose space and time-dependent transition probabilities in each step are extracted from the amplitudes of the DTQW. This approach yields a non-homogeneous CRW whose probability distribution replicates that of the corresponding DTQW. Although effective in reproducing certain statistics, these methods necessarily break the spatial and temporal structure of the DTQW and demand access to quantum amplitudes at each step, an approach analogous to Bohmian mechanics[9], where the quantum wave function guides particle trajectories.
@@
In contrast, alternative formulations of quantum mechanics—akin to the hydrodynamic approaches of Holland and Poirier [10, 11, 12, 13, 14, 15] —eschew direct reliance on the wave function, instead constructing real-valued fields or potentials that reproduce quantum evolution without amplitude-based guidance. Inspired by this distinction, we propose an interacting particles framework for modeling DTQWs that never invokes complex amplitudes. All update rules are real‑valued processes whose parameters depend only on the parameters of Quantum walker’s coin and shift operations. Crucially, our model mirrors the structural properties of the original quantum walk system. When the Quantum walk has uniform coin and shift operations, the classical interacting particle dynamics remains homogeneous in space and time, employing identical update rules at each node and time step. If Quantum walker’s coin or shift operations vary across space or time, the corresponding interacting particle model adapts its parameters in the same pattern, preserving the structural variations of Quantum walk. The Classical interacting particle model yields two main advantages. First, preserving the structural variations of the Quantum walk simplifies the analysis, while allowing our method to be naturally extended to higher-dimensional graphs. Second, by avoiding amplitude-based guidance, the classical model brings into focus a fundamental probabilistic mechanism underlying quantum walk spreading and interference, insights that may carry foundational significance.
@@
[…]
@@
Here $\mathbb{I}_{p}$ is the identity operator on $\mathcal{H}_{c}$. The process of $W$ is iterated without resorting to the intermediate measurements to realize a large number of steps of the quantum walk.
@@
is sufficient to describe the most general form of the discrete-time quantum walk, where $\xi,\theta,\zeta\in\mathbb{R}$ and $\theta\in[0,\pi/2]$. By choosing different $(\xi,\theta,\zeta)$, one obtains a variety of discrete-time walks with distinct dynamical features. For example one can write $H=C(0,\frac{\pi}{4},0)$.
@@
Suppose
$\psi_{0}^{\prime}=r_{0}^{\prime}e^{i\phi_{0}^{\prime}},\qquad\psi_{1}^{\prime}=r_{1}^{\prime}e^{i\phi_{1}^{\prime}},$
then one may write
$r_{0}^{\prime}e^{i\phi_{0}^{\prime}}$ $=a_{0}^{\prime}+ib_{0}^{\prime},$ (10)
$r_{1}^{\prime}e^{i\phi_{1}^{\prime}}$ $=a_{1}^{\prime}+ib_{1}^{\prime},$ (11)
where
$a_{0}^{\prime}$ $=\tfrac{1}{\sqrt{2}}\!\left(r_{0}\cos\phi_{0}+r_{1}\cos\phi_{1}\right),$
$b_{0}^{\prime}$ $=\tfrac{1}{\sqrt{2}}\!\left(r_{0}\sin\phi_{0}+r_{1}\sin\phi_{1}\right),$
$a_{1}^{\prime}$ $=\tfrac{1}{\sqrt{2}}\!\left(r_{0}\cos\phi_{0}-r_{1}\cos\phi_{1}\right),$
$b_{1}^{\prime}$ $=\tfrac{1}{\sqrt{2}}\!\left(r_{0}\sin\phi_{0}-r_{1}\sin\phi_{1}\right).$ (12)
The corresponding probabilities of measuring the transformed state in the computational basis are
$P_{0}^{\prime}$ $={r_{0}^{\prime}}^{2}=\tfrac{1}{2}\!\left(r_{0}^{2}+r_{1}^{2}+2r_{0}r_{1}\cos(\phi_{0}-\phi_{1})\right),$ (13)
$P_{1}^{\prime}$ $={r_{1}^{\prime}}^{2}=\tfrac{1}{2}\!\left(r_{0}^{2}+r_{1}^{2}-2r_{0}r_{1}\cos(\phi_{0}-\phi_{1})\right).$ (14)
The phases $\phi_{0}^{\prime}$ and $\phi_{1}^{\prime}$ are obtained from
$\phi_{0}^{\prime}$ $=\mathrm{atan2}\left(b_{0}^{\prime}/a_{0}^{\prime}\right),$ $\phi_{1}^{\prime}$ $=\mathrm{atan2}\left(b_{1}^{\prime}/a_{1}^{\prime}\right).$ (15)
Figure 1: Schematic illustration of the system of two-box. A collection of $N$ balls is distributed between two boxes labeled 0 and $1$, each carrying an associated phase tag $\eta_{0}$ and $\eta_{1}$. The transformation process $\mathbb{T}_{c}$ updates both the occupation numbers and the phase tags, mapping the initial configuration to a new one according to prescribed rules.
Now consider a system of two boxes (Fig. 1), labeled 0 and $1$, and a total collection of $N$ particles, hereafter referred to as balls. The various operations on this system are categorized into three classes: preparation ($\mathbb{P}$), transformation ($\mathbb{T}_{c}$), and measurement ($\mathbb{M}$). Throughout the protocol, these operations are implemented by an experimenter, which may be either a human agent or an automated device, and whose role is to apply the prescribed rules to the system without introducing any additional dynamics beyond those specified by the model.
@@
In a typical preparation procedure $\mathbb{P}$, the experimenter begins by selecting one ball uniformly at random from the collection of $N$ balls and marking it for future reference. The full set of $N$ balls (including the marked one) is then distributed between the two boxes. This allocation proceeds in one of two ways, each chosen with equal probability:
@@
  • Either $N_{0}=\lfloor W_{N}\rfloor$ balls are randomly selected from the collection and placed into box 0, with the remaining $N_{1}=N-N_{0}$ balls assigned to box $1$;
@@
  • Or $N_{0}=\lceil W_{N}\rceil$ balls are randomly selected from the collection and placed into box 0, with the remaining $N_{1}=N-N_{0}$ balls placed into box $1$.
Here, $\lfloor\cdot\rfloor$ and $\lceil\cdot\rceil$ denote the floor and ceiling functions, respectively and $W_{N}=\rho_{0}N$. The parameter $\rho_{0}\in[0,1]$ controls the expected proportion of balls assigned to box 0. Throughout the selection procedure, all balls—including the marked one—are treated as identical.
In addition, each box $k$ (with $k=0,1$) is assigned a real-valued phase tag $\eta_{k}$. The resulting prepared state is denoted by $\textit{B}^{N_{0},N_{1}}_{\eta_{0},\eta_{1}}$.
We now define the transformation procedure $\mathbb{T}_{c}$ acting on the state $\textit{B}^{N_{0},N_{1}}_{\eta_{0},\eta_{1}}$. In this procedure, the experimenter first introduces a function
$\tilde{N}_{0}=\tilde{N}_{0}(N_{0},N_{1},\eta_{0},\eta_{1}),$
and then assigns the updated occupation number $N_{0}^{\prime}$ by choosing either $\lfloor\tilde{N}_{0}\rfloor$ or $\lceil\tilde{N}_{0}\rceil$ with equal probability.
The system is subsequently updated according to the following rule:
  • If $N_{0}^{\prime}>N_{0}$, the experimenter randomly selects $N_{0}^{\prime}-N_{0}$ balls from box $1$ and transfers them to box 0.
  • If $N_{0}^{\prime}<N_{0}$, the experimenter randomly selects $N_{0}-N_{0}^{\prime}$ balls from box 0 and transfers them to box $1$.
The phase tags $\eta_{0}$ and $\eta_{1}$ are also updated to new values $\eta_{0}^{\prime}$ and $\eta_{1}^{\prime}$, where
$\eta_{k}^{\prime}=\eta_{k}^{\prime}(N_{0},N_{1},\eta_{0},\eta_{1}),\qquad k=0,1.$
Putting everything together, the transformation procedure is summarized as
$\mathbb{T}_{c}:\;\textit{B}^{N_{0},N_{1}}_{\eta_{0},\eta_{1}}\;\longrightarrow\;\textit{B}^{N_{0}^{\prime},N_{1}^{\prime}}_{\eta_{0}^{\prime},\eta_{1}^{\prime}}.$
@@
The measurement procedure, denoted by $\mathbb{M}$, is implemented as follows. In this process, the experimenter places all balls into the box that contains the marked ball, and the measurement outcome is identified with the label of that box. Both phase tags are then reset to zero.
Accordingly, the post-measurement state becomes
  • $\textit{B}^{N,0}_{0,0}$ if the marked ball is in box 0;
  • $\textit{B}^{0,N}_{0,0}$ if the marked ball is in box $1$.
Suppose the experimenter prepares an initial state $\textit{B}^{N_{0},N_{1}}_{\phi_{0},\phi_{1}}$, with the identifications $\eta_{0}=\phi_{0}$ and $\eta_{1}=\phi_{1}$. Here $\phi_{0}$ and $\phi_{1}$ are the phases of the complex probability amplitudes $\psi_{0}$ and $\psi_{1}$ defined in Eq. (5). The occupation numbers $N_{0}$ and $N_{1}$ are obtained through the preparation procedure $\mathbb{P}$ by choosing $\rho_{0}=r_{0}^{2}$, so that $W_{N}=r_{0}^{2}N$, where $r_{0}^{2}$ is the probability of obtaining the outcome $\lvert 0\rangle$ from the quantum state in Eq. (5).
If the experimenter then applies the measurement procedure $\mathbb{M}$, the probability of obtaining outcome 0 in the limit of large $N$ is
@@
$\lim_{N\to\infty}\frac{\langle N_{0}\rangle}{N}=\lim_{N\to\infty}\frac{\lfloor{N}_{0}\rfloor+\lceil{N}_{0}\rceil}{2N}=r_{0}^{2},$ (16)
while the probability of obtaining outcome $1$ becomes
$\lim_{N\to\infty}\left(1-\frac{\langle N_{0}\rangle}{N}\right)=r_{1}^{2}.$ (17)
Now suppose the experimenter starts from the same initial state $\textit{B}^{N_{0},N_{1}}_{\phi_{0},\phi_{1}}$ and applies a specific transformation to this state. This transformation is designed to reproduce the action of the Hadamard coin (Eqs. (13), (14), and (15)) within the present classical box and balls framework.
In this $\mathbb{T}_{c}$ procedure, the experimenter follows the transformation scheme defined below. First, the quantity
$\tilde{N}_{0}=\frac{N_{0}+N_{1}}{2}+\sqrt{N_{0}N_{1}}\,\cos(\phi_{1}-\phi_{0})$ (18)
is introduced, together with the auxiliary functions
$B_{k}$ $=\sqrt{\frac{N_{0}}{N}}\,\sin\phi_{0}+(-1)^{k}\sqrt{\frac{N_{1}}{N}}\,\sin\phi_{1},$ (19)
$A_{k}$ $=\sqrt{\frac{N_{0}}{N}}\,\cos\phi_{0}+(-1)^{k}\sqrt{\frac{N_{1}}{N}}\,\cos\phi_{1}.$ (20)
The updated phase tags associated with the boxes are then determined by
$\eta_{k}^{\prime}(N_{0},N,\phi_{0},\phi_{1})=\mathrm{atan2}\left(\frac{B_{k}}{A_{k}}\right),$ (21)
and the experimenter sets the updated number of balls in box 0, denoted $N_{0}^{\prime}$, to either
@@
$\left\lfloor\frac{N}{2}+\sqrt{N_{0}N_{1}}\cos(\phi_{1}-\phi_{0})\right\rfloor\quad\text{or}\quad\left\lceil\frac{N}{2}+\sqrt{N_{0}N_{1}}\cos(\phi_{1}-\phi_{0})\right\rceil,$
each choice occurring with equal probability. The method by which the experimenter obtained this updated state has been described by the general method of the transformation process $\mathbb{T}_{c}$ introduced previously.
@@
The specific forms of $\tilde{N}_{0}$, $A_{k}$, and $B_{k}$ in the transformation $\mathbb{T}_{H}$ are chosen such that, in the limit $N\to\infty$, the probability of obtaining the measurement outcome 0 satisfies
@@
while the probability of outcome $1$ becomes
@@
In addition, the phase-tag updates are also satisfy
$\lim_{N\to\infty}\langle\eta_{k}^{\prime}(N_{0},N,\phi_{0},\phi_{1})\rangle={\phi_{k}}^{\prime}.$ (24)
As $N\to\infty$, the measurement-outcome probabilities generated by this classical procedure converge to those produced by the Hadamard transformation on the quantum state of Eq. (5) .
For finite $N$, deviations are unavoidable; however, by taking $N$ sufficiently large, such quantum operations can be approximated with arbitrary precision within the present classical framework.
Moreover, this method can be generalized to simulate other single-qubit quantum coin operations, as expressed in Eq. (4). One can obtain the generalized forms of $\tilde{N}_{0}$ and $\eta_{i}^{\prime}$ in the corresponding transformation process as follows :
[…]
$\eta_{k}^{\prime}=\mathrm{atan2}\left(\frac{\mathbb{B}_{k}}{\mathbb{A}_{k}}\right),$ (28)
where,
$\mathbb{B}_{k}=\sqrt{\frac{N_{0}}{N}}\cos\theta\sin(\xi+\phi_{0})+(-1)^{k}\sqrt{\frac{N_{1}}{N}}\sin\theta\sin(\zeta+\phi_{1}),$
$\mathbb{A}_{k}=\sqrt{\frac{N_{0}}{N}}\cos\theta\cos(\xi+\phi_{0})+(-1)^{k}\sqrt{\frac{N_{1}}{N}}\sin\theta\cos(\zeta+\phi_{1}).$
Finally, we introduce the boxes-and-balls model corresponding to a 1D discrete-time quantum walk (DTQW) with a generalized $SU(2)$ coin [Eq. (4)], and an initial state of the form:
Figure 2: Schematic illustration of the boxes-and-balls model for a discrete-time quantum walk. Each lattice site $x\in\mathbb{Z}$ contains two boxes labeled by $c=0,1$, with associated phase tags $\eta_{x,c}$. At time $t=0$, the $N$ balls (one of which is marked) are distributed according to the preparation process $\mathbb{P}$. The transformation $\mathbb{T}_{c}$ updates both the occupation numbers $N_{0,c}$ and the phase tags, followed by the conditional shift that moves the balls and associated phase tags to left or right depending on $c$. The configuration at $t=1$ illustrates the resulting redistribution after one full step.
Figure 3: Probability distributions arising from the $\mathbb{M}$ process in the boxes-and-balls model of the Hadamard coined quantum walk, with initial state $\frac{1}{\sqrt{2}}\lvert 0\rangle-\frac{i}{\sqrt{2}}\lvert 1\rangle$ and at time $t=100$, are shown for different total numbers of particles NN. As NN increases, the distribution approaches that of the ideal Hadamard quantum walk, indicated by the dotted line.
@@
In this setting, the dynamics is implemented through the three processes $\mathbb{P}$, $\mathbb{T}_{c}$, and $\mathbb{M}$ within a generalized framework. To each lattice site $x\in\mathbb{Z}$ we assign two boxes and a phase tag to each box, labeled $(x,c)$ and $\eta_{x,c}$ respectively, with $c\in\{0,1\}$ (Fig. 2). We begin with a collection of $N$ balls, one of which is randomly selected and marked by the experimenter. The full set of $N$ balls is then distributed among the boxes according to a preparation protocol that will be specified later.
At any time $t$, the state of the system is completely characterized by the nonnegative integers
$N_{x,c}(t)=\text{number of balls in box }(x,c),\qquad c\in\{0,1\},$
together with a phase tag $\eta_{x,c}(t)\in[0,2\pi)$ associated with each box $(x,c)$. The occupation numbers $N_{x,c}$ satisfy the global constraint
@@
At time $t=0$, only the two boxes at the origin contains the balls ; all others are empty:
@@
Within the origin boxes, we set